OpenStudy (anonymous):

Can you please explain the change from i dt = C dv to the equivilant equation integral from v(t_0) to v(t) of dx = 1/C integral from t_0 to t of i dτ

OpenStudy (anonymous):

\[i* dt = C* dv , or \int\limits_{v(t_o)}^{v(t)} dx =1/C \int\limits_{t_0}^{t}idτ\]